# Optimal Low-Depth Quantum Signal-Processing Phase Estimation

### v1

@@ F.4 Comparing with quantum Cramer-Rao lower bound
where Lemma 20 is used. Note that the upper bound on the QFI derived here is independent of the initialization $\varrho$, and the formalism of QFI provides a bound on the inference variance regardless of the choice of measurements. Hence, we achieve a lower bound on the variance of $\hat{\theta}$ with respect to any initialization, measurement, and classical data processing by invoking the quantum Cramer-Rao bound:
@@
    The importance of the phase matching condition is also reflected in the analysis of QFI. It is worth noting that when we set $\omega_{j}=\varphi$, the QFI of a single experiment is $\mathfrak{F}_{j}(\theta)|_{\omega_{j}=\varphi}\leq 4U_{d-1}^{2}(1)=4d^{2}$ which attains the maximum of the Chebyshev polynomial. However, in practice, due to the absence of accurate information of $\varphi$, our method samples the data on equally spaced $\omega$-points and processes the data via FFT to isolate the dependencies of $\theta$ and $\varphi$. This further sampling procedure averages the QFI and lowers its value from $4d^{2}$ to $4d$.
@@
Appendix G Analysis of realistic error
@@
G.1 Depolarizing error
@@
We numerically test the accuracy of these estimators in Section G.3. In Table 1, we list the effective depolarizing error rate on the single-excitation subspace inferred from the exponential decay of circuit fidelities derived from QSPE methods on our experimental measurements. These values agree with our estimation results using a conventional cross-entropy benchmark that is much slower to run and requires a $10X$ deeper circuit for randomization.
@@
Table 1: Qubit pairs and the inferred effective error rate on the single-excitation subspace. The error rate is estimated by the regression with respect to the exponential decay. The regression data are the circuit fidelity estimated from QSPE in Figure 7 (top-right panel).
G.2 Time-dependent error
@@
G.3 Numerical performance of the estimation against depolarizing error and time-dependent drift error
@@
where $r$ is the error rate. At the same time, the quantum circuit is subject to drift error according to Equations 169 and 170.
@@
Figure 16: Estimating circuit fidelity using QSPE. The reference value $\alpha_{\mathrm{DEM}}$ is the circuit fidelity estimated from the digital error model. The sources of noise in the numerical experiments are Monte Carlo sampling error, depolarizing error and drift error. The depolarizing error rate is set to $r=1\times 10^{-3}$ and the number of measurement samples is set to $M=1\times 10^{5}$. The parameters of FsimGate are set to $\theta=1\times 10^{-3}$ and $\chi=5\pi/32$. The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
@@
In Figures 17 and 18, we test our proposed metrology scheme in the presence of Monte Carlo sampling error, depolarizing error, and time-dependent error. Although the system is subjected to realistic errors, the numerical results suggest that the QSPE estimators show some robustness against errors, and they can give reasonable estimation results with one or two correct digits. Furthermore, the accuracy of $\varphi$-estimation is also not fully contaminated by the time-dependent error on it. The improvement due to the peak fitting becomes less significant under realistic errors because the structure of the highest peak is heavily distorted in the presence of realistic errors. More interestingly, the numerical results show the accuracy of $\theta$-estimation does not decay and even increases after some $d^{*}$. This transition is due to a tradeoff. When $d$ becomes larger, the inference is expected to be more accurate because the gate parameters are more amplified. However, in the presence of realistic error, the FsimGate is subjected to both time-independent errors and time-dependent drift errors. A quantum circuit with more FsimGate s violates the model derived from the noiseless setting more. The competition between these two opposite effects makes the estimation error attain some minimum at $d^{*}$. This observation also suggests that in the experimental deployment, one can consider using a moderate $d$ with respect to the tradeoff.
@@
In Figure 18, we perform the numerical simulation with variable swap angle and number of measurement samples. Similar to the case of Monte Carlo sampling error, the estimation results are less accurate when $\theta$ is small because of the insufficient SNR. The numerical results indicate that the estimation accuracy cannot be further improved after the number of measurement samples is greater than some $M^{*}$. That is because increasing $M$ can only mitigate Monte Carlo sampling error. When $M$ is large enough, the sources of errors are dominated by depolarizing error and time-dependent drift error, which cannot be sufficiently mitigated by large $M$. Combing with the discussion on $d^{*}$, the numerical results suggest that the experimental deployment does not require an extremely large $d$ and $M$, and using a moderate choice of $d^{*}$ and $M^{*}$ suffices to get some accurate estimation.
@@
Figure 17: Accuracy of estimators as a function of the number of FsimGate s. The sources of noise in the numerical experiments are Monte Carlo sampling error, depolarizing error and time-dependent drift error. The depolarizing error rate is set to $r=1\times 10^{-3}$ and the number of measurement samples is set to $M=1\times 10^{5}$. The swap angle is set to $\theta=1\times 10^{-3}$ and the phase parameter is set to $\chi=5\pi/32$. The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
@@
Figure 18: Accuracy of estimators as a function of the number of measurement samples. The sources of noise in the numerical experiments are Monte Carlo sampling error, depolarizing error and time-dependent drift error. The depolarizing error rate is set to $r=1\times 10^{-3}$. The circuit degree is set to $d=50$ and the FsimGate phase parameter is set to $\chi=5\pi/32$.The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
G.4 Readout error
@@
where $\mathbb{P}(\mathrm{binary}(j)|\mathrm{binary}(j))$ is the conditional probability of measuring the qubits with the bit-string $\mathrm{binary}(j)$ given that the quantum state is |$\ket{\mathrm{binary}(i)}$. The sum of each row of the confusion matrix is equal to one due to the normalization of probability. The confusion matrix can be determined by performing additional quantum experiments in which $I\otimes I$, $I\otimes X$, $X\otimes I$, and $X\otimes X$ are measured to determine each row, respectively. If the probability vector from the measurement with readout error is $\vec{q}^{\mathrm{exp}}=(q^{\mathrm{exp}}(00),q^{\mathrm{exp}}(01),q^{\mathrm{exp}}(10),q^{\mathrm{exp}}(11))^{\top}$, the probability vector after correcting the readout error is given by inverting the confusion matrix
$\vec{p}^{\mathrm{exp}}=(p^{\mathrm{exp}}(00),p^{\mathrm{exp}}(01),p^{\mathrm{exp}}(10),p^{\mathrm{exp}}(11))^{\top}=\left(R^{\top}\right)^{-1}\vec{q}^{\mathrm{exp}}.$ (174)
@@
Let $\vec{p}^{\mathrm{exp}}_{\mathrm{fs}}$ be the probability vector computed by inverting the confusion matrix estimated by finite samples. To achieve the bounded error $\norm{\vec{p}^{\mathrm{exp}}-\vec{p}^{\mathrm{exp}}_{\mathrm{fs}}}_{2}\leq\epsilon$ with confidence level $1-\alpha$, it suffices to set the number of measurement samples in each experiment determining the confusion matrix as
@@
is therefore an estimate to the measurement probability since $\mathbb{E}\left(\mathbf{q}^{(u)}\right)=\mathbf{p}^{(u)}$. However, the statistical fluctuation makes the estimate deviates the exact probability. Applying Hoeffding’s inequality, we have
@@
$\norm{\vec{p}^{\mathrm{exp}}-\vec{p}^{\mathrm{exp}}_{\mathrm{fs}}}_{2}\leq\norm{R_{\mathrm{fs}}^{-1}-R^{-1}}_{2}\norm{\vec{q}^{\mathrm{exp}}}_{2}\leq\norm{R_{\mathrm{fs}}^{-1}-R^{-1}}_{2}\norm{\vec{q}^{\mathrm{exp}}}_{1}\leq\frac{\widetilde{\epsilon}\kappa^{2}}{1-\widetilde{\epsilon}\kappa}.$ (187)
@@
Thus, to achieve the bounded error $\norm{\vec{p}^{\mathrm{exp}}-\vec{p}^{\mathrm{exp}}_{\mathrm{fs}}}_{2}\leq\epsilon$ with confidence level $1-\alpha$, it suffices to set the number of measurement samples in each experiment determining the confusion matrix as
@@
Appendix H Computing the polynomial representation on a special set of points
@@
Furthermore, $Q_{\omega}^{(d)}$ and $\mathrm{Im}\left(e^{-\mathrm{i}\omega}P_{\omega}^{(d)}\right)$ can be determined from the recurrence relation in Equations 191 and 192
@@
Appendix I Analysis of periodic calibration: variance lower bound and shortcomings
I.1 Overview of the methodology of periodic calibration
In this subsection, we provide an overview of periodic calibration, also known as Floquet calibration, which is proposed in [2, Appendix C] and [23, Appendix A]. Periodic calibration is a generalization of the robust single-qubit gate calibration [14] to entangling gates. The main component of the quantum circuit for performing periodic calibration is equivalent to the periodic part used in our method (see the shaded part in Figure 1). However, instead of the initialization in terms of Bell states, periodic calibration measures the transition probability between tensor product states |$\ket{01}$ and |$\ket{10}$, namely, between logical quantum state |$\ket{0_{\ell}}$ and |$\ket{1_{\ell}}$. The exact parametric expression of this probability can be derived from the results presented in Appendix C, which is consistent with that in literature (upon convention difference):
@@ I.1 Overview of the methodology of periodic calibration
I.2 Optimality analysis using Fisher information and Cramér-Rao bound
@@
As we derived in Equation 157 in Section F.4, the quantum Fisher information that a quantum circuit using $\ell$ $U$-gates can maximally contribute is $\mathfrak{F}_{\max}=4M\ell^{2}$. Hence, the coincidence between the classical Fisher information with phase-matching condition and the maximal quantum Fisher information indicates that the optimality is attained at phase-matching because the Fisher information attains its maximum.
@@
Figure 19: Cramér-Rao lower bound (CRLB) on the estimation variance of $\theta$ using periodic calibration. The number of measurement samples is set to $M=1\times 10^{5}$.
I.3 Violation of phase-matching condition implies exponentially worse estimation variance
To understand the importance of the phase-matching condition and the plateau of estimation variance, we provide an approximation to the Fisher information when the swap angle $\theta$ is small. For simplicity, we focus on the quantum circuit with a fixed depth parameter $d$ rather than a series of circuits with variable depths. It is referred to as $J_{\theta}$ below. We remark that the focus on the fixed-depth Fisher information does not loose generality due to the additivity of Fisher information.
@@ I.3 Violation of phase-matching condition implies exponentially worse estimation variance
Figure 20: Approximation analysis of Cramér-Rao lower bound (CRLB). (a) Fisher information of a fixed depth $J_{\theta}$ (blue solid line) and its approximation in Equation 204 (orange dashed line). The shaded area is the principal peak with |$\absolutevalue{\varphi}\leq\pi/d$. We set $d=10$ and $M=1$. (b) CRLB (scatters) and its approximation (dashed lines) based on Equation 204. We set $M=1\times 10^{5}$.
@@
I.4 Practical challenges due to the complex optimization landscape
@@
Figure 21: Landscape analysis of periodic calibration. We set the exact value $\theta=1$ for all figures. (a) Optimization landscape of different $\varphi$ angles. We set the maximum depth $d=2^{1}0=1024$ and consider an ideal case where infinite many measurement samples are used $M=\infty$. (b) Optimization landscape of different maximum depth parameters $d$. We consider an ideal case where infinite many measurement samples are used $M=\infty$, and the phase-matching condition is satisfied $\varphi=0$. (c) Optimization landscape of different numbers of measurement samples $M$. We set $d=2^{8}=256$. We consider an ideal case where the phase-matching condition is satisfied $\varphi=0$.

### v2

@@ Acknowledgments
where Lemma 18 is used. Note that the upper bound on the QFI derived here is independent of the initialization $\varrho$, and the formalism of QFI provides a bound on the inference variance regardless of the choice of measurements. Hence, we achieve a lower bound on the variance of $\hat{\theta}$ with respect to any initialization, measurement, and classical data processing by invoking the quantum Cramér-Rao bound:
@@
    The importance of the phase matching condition is also reflected in the analysis of QFI. It is worth noting that when we set $\omega_{j}=\varphi$, the QFI of a single experiment is $\mathfrak{F}_{j}(\theta)|_{\omega_{j}=\varphi}\leq 4U_{d-1}^{2}(1)=4d^{2}$ which attains the maximum of the Chebyshev polynomial. However, in practice, due to the absence of accurate information of $\varphi$, our method samples the data on equally spaced $\omega$ points and processes the data via FFT to isolate the dependencies of $\theta$ and $\varphi$. This further sampling procedure averages the QFI and lowers its value from $4d^{2}$ to $4d$.
@@
Supplementary Information 7 Analysis of realistic error
@@
Supplementary Information 7.1 Depolarizing error
@@
We numerically test the accuracy of these estimators in Section Supplementary Information 7.3. In Table 1, we list the effective depolarizing error rate on the single-excitation subspace inferred from the exponential decay of circuit fidelities derived from QSPE methods on our experimental measurements. These values agree with our estimation results using a conventional cross-entropy benchmark that is much slower to run and requires a $10X$ deeper circuit for randomization.
@@
Supplementary Table 1: Qubit pairs and the inferred effective error rate on the single-excitation subspace. The error rate is estimated by the regression with respect to the exponential decay. The regression data are the circuit fidelity estimated from QSPE in Figure 4 in the main text (top-right panel).
Supplementary Information 7.2 Time-dependent error
@@
Supplementary Information 7.3 Numerical performance of the estimation against depolarizing error and time-dependent drift error
@@
where $r$ is the error rate. At the same time, the quantum circuit is subject to drift error according to Equations 157 and 158.
@@
Supplementary Figure 11: Estimating circuit fidelity using QSPE. The reference value $\alpha_{\mathrm{DEM}}$ is the circuit fidelity estimated from the digital error model. The sources of noise in the numerical experiments are Monte Carlo sampling error, depolarizing error and drift error. The depolarizing error rate is set to $r=1\times 10^{-3}$ and the number of measurement samples is set to $M=1\times 10^{5}$. The parameters of FsimGate are set to $\theta=1\times 10^{-3}$ and $\chi=5\pi/32$. The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
@@
In Figures 12 and 13, we test our proposed metrology scheme in the presence of Monte Carlo sampling error, depolarizing error, and time-dependent error. Although the system is subjected to realistic errors, the numerical results suggest that the QSPE estimators show some robustness against errors, and they can give reasonable estimation results with one or two correct digits. Furthermore, the accuracy of $\varphi$-estimation is also not fully contaminated by the time-dependent error on it. The improvement due to the peak fitting becomes less significant under realistic errors because the structure of the highest peak is heavily distorted in the presence of realistic errors. More interestingly, the numerical results show the accuracy of $\theta$-estimation does not decay and even increases after some $d^{*}$. This transition is due to a tradeoff. When $d$ becomes larger, the inference is expected to be more accurate because the gate parameters are more amplified. However, in the presence of realistic error, the FsimGate is subjected to both time-independent errors and time-dependent drift errors. A quantum circuit with more FsimGate s violates the model derived from the noiseless setting more. The competition between these two opposite effects makes the estimation error attain some minimum at $d^{*}$. This observation also suggests that in the experimental deployment, one can consider using a moderate $d$ with respect to the tradeoff.
@@
In Figure 13, we perform the numerical simulation with variable swap angle and number of measurement samples. Similar to the case of Monte Carlo sampling error, the estimation results are less accurate when $\theta$ is small because of the insufficient SNR. The numerical results indicate that the estimation accuracy cannot be further improved after the number of measurement samples is greater than some $M^{*}$. That is because increasing $M$ can only mitigate Monte Carlo sampling error. When $M$ is large enough, the sources of errors are dominated by depolarizing error and time-dependent drift error, which cannot be sufficiently mitigated by large $M$. Combing with the discussion on $d^{*}$, the numerical results suggest that the experimental deployment does not require large $d$ and $M$, and using a moderate choice of $d^{*}$ and $M^{*}$ suffices to get some accurate estimation.
@@
Supplementary Figure 12: Accuracy of estimators as a function of the number of FsimGate s. The sources of noise in the numerical experiments are Monte Carlo sampling error, depolarizing error and time-dependent drift error. The depolarizing error rate is set to $r=1\times 10^{-3}$ and the number of measurement samples is set to $M=1\times 10^{5}$. The swap angle is set to $\theta=1\times 10^{-3}$ and the phase parameter is set to $\chi=5\pi/32$. The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
@@
Supplementary Figure 13: Accuracy of estimators as a function of the number of measurement samples. The sources of noise in the numerical experiments are Monte Carlo sampling error, depolarizing error and time-dependent drift error. The depolarizing error rate is set to $r=1\times 10^{-3}$. The circuit degree is set to $d=50$ and the FsimGate phase parameter is set to $\chi=5\pi/32$.The error bar of each point stands for the confidence interval derived from $96$ independent repetitions.
Supplementary Information 7.4 Readout error
@@
where $\mathbb{P}(\mathrm{binary}(j)|\mathrm{binary}(j))$ is the conditional probability of measuring the qubits with the bit-string $\mathrm{binary}(j)$ given that the quantum state is |$\ket{\mathrm{binary}(i)}$. The sum of each row of the confusion matrix is equal to one due to the normalization of probability. The confusion matrix can be determined by performing additional quantum experiments in which $I\otimes I$, $I\otimes X$, $X\otimes I$, and $X\otimes X$ are measured to determine each row, respectively. If the probability vector from the measurement with readout error is $\mathbf{q}^{\mathrm{exp}}=(q^{\mathrm{exp}}(00),q^{\mathrm{exp}}(01),q^{\mathrm{exp}}(10),q^{\mathrm{exp}}(11))^{\top}$, the probability vector after correcting the readout error is given by inverting the confusion matrix
$\mathbf{p}^{\mathrm{exp}}=(p^{\mathrm{exp}}(00),p^{\mathrm{exp}}(01),p^{\mathrm{exp}}(10),p^{\mathrm{exp}}(11))^{\top}=\left(R^{\top}\right)^{-1}\mathbf{q}^{\mathrm{exp}}.$ (162)
@@
Let $\mathbf{p}^{\mathrm{exp}}_{\mathrm{fs}}$ be the probability vector computed by inverting the confusion matrix estimated by finite samples. To achieve the bounded error $\norm{\mathbf{p}^{\mathrm{exp}}-\mathbf{p}^{\mathrm{exp}}_{\mathrm{fs}}}_{2}\leq\epsilon$ with confidence level $1-\alpha$, it suffices to set the number of measurement samples in each experiment determining the confusion matrix as
@@
is therefore an estimate to the measurement probability since $\mathbb{E}\left(\mathbf{q}^{(u)}\right)=\mathbf{p}^{(u)}$. However, the statistical fluctuation makes the estimate deviate the exact probability. Applying Hoeffding’s inequality, we have
@@
$\norm{\mathbf{p}^{\mathrm{exp}}-\mathbf{p}^{\mathrm{exp}}_{\mathrm{fs}}}_{2}\leq\norm{R_{\mathrm{fs}}^{-1}-R^{-1}}_{2}\norm{\mathbf{q}^{\mathrm{exp}}}_{2}\leq\norm{R_{\mathrm{fs}}^{-1}-R^{-1}}_{2}\norm{\mathbf{q}^{\mathrm{exp}}}_{1}\leq\frac{\widetilde{\epsilon}\kappa^{2}}{1-\widetilde{\epsilon}\kappa}.$ (175)
@@
Thus, to achieve the bounded error $\norm{\mathbf{p}^{\mathrm{exp}}-\mathbf{p}^{\mathrm{exp}}_{\mathrm{fs}}}_{2}\leq\epsilon$ with confidence level $1-\alpha$, it suffices to set the number of measurement samples in each experiment determining the confusion matrix as
@@
Supplementary Information 7.5 Initial state preparation error
In this subsection, we analyze the effect of the error in the initialization of Bell states in the quantum circuits. One case is that the error brings the initial state out of the subspace $\mathrm{span}\{\ket{0_{\ell}},\ket{1_{\ell}}\}$. Because the rest of the circuit preserves that two-dimensional subspace, such error can be mitigated by post-selecting the measurement outcomes in the two-dimensional subspace. Hence, in the rest of this subsection, we focus on the initial state error within the two-dimensional subspace.
[…]
$\begin{split}p_{\beta}^{\prime}&=\absolutevalue{\bra{0_{\ell}}\mathcal{U}^{(d)}(\omega;\theta,\varphi,\chi)\ket{\beta^{\prime}}}^{2}\approx\absolutevalue{\bra{0_{\ell}}\mathcal{U}^{(d)}(\omega;\theta,\varphi,\chi)\ket{\beta}-\mathrm{i}\eta\bra{0_{\ell}}\mathcal{U}^{(d)}(\omega;\theta,\varphi,\chi)K\ket{\beta}}^{2}+\mathcal{O}(\eta^{2})\\ &\approx p_{\beta}\left(1+\mathcal{O}(p_{\beta}^{-1/2}\eta\bra{0_{\ell}}\mathcal{U}^{(d)}(\omega;\theta,\varphi,\chi)K\ket{\beta})\right)+\mathcal{O}(\eta^{2})\\ &\approx p_{\beta}+\mathcal{O}(\eta\bra{0_{\ell}}\mathcal{U}^{(d)}(\omega;\theta,\varphi,\chi)K\ket{\beta})+\mathcal{O}(\eta^{2}+d\theta\eta)\end{split}$ (178)
Here, the second and the third lines use the conclusion that $\bra{0_{\ell}}\mathcal{U}^{(d)}(\omega;\theta,\varphi,\chi)\ket{\beta}=\mathcal{O}(\sqrt{p_{\beta}})=\mathcal{O}(1)+\mathcal{O}(d\theta)$ when $d\theta\ll 1$. Note that Equations 140 and 142 indicate that
$\mathcal{O}(\mathcal{U}^{(d)}(\omega;\theta,\varphi,\chi))=\mathcal{O}(I+\mathrm{i}\theta XU_{d-1}(\cos(\omega-\varphi)))+\mathcal{O}((d\theta)^{2}).$ (179)
Note that the identity matrix in the expansion only contributes to the constant term, which we explicitly annotate in the following result:
$p_{\beta}^{\prime}\approx p_{\beta}+\text{const}\times\eta+\text{const}\times\eta\theta U_{d-1}(\omega-\varphi)+\mathcal{O}(\eta^{2}+d\theta\eta).$ (180)
We remark that the constant-shift term $\text{const}\times\eta$ only affects the zeroth Fourier mode, which can be dropped by post-selecting Fourier modes of positive indices. This perturbative analysis implies that when subjected to initial state error with magnitude $\eta$, each Fourier component has an additive error magnitude $\mathcal{O}(\eta\theta)$. Thus, the initial state error contributes to a relative error in the estimation, $|\hat{\theta}-\theta|/\theta=\mathcal{O}(\eta)$.
The perturbative analysis above gives intuition in understanding the initial state error. However, the derivation relies on several approximations that may not cover all cases. In the following text, we perform numerical analysis to have a more comprehensive understanding of the estimation error induced by initial state error. We turn off all other noises including sampling error to focus on the effect of initial state error. We perform three different strategies to add initial state error to the system.
  1. 1.
    “identical”: The initial state error $E_{\eta}=\exp(-\mathrm{i}\eta(X+Y)/\sqrt{2})$ is identical in all experiments for any $\beta\in\{+,\mathrm{i}\}$ and any $\omega$.
  2. 2.
    “fixed”: The initial state error $E_{\eta}$ only depends on the initial state to be prepared. It is fixed in all experiments for any $\omega$ but takes two possible values $E_{\eta,+}=\exp(-\mathrm{i}\eta(X+Y)/\sqrt{2})$ and $E_{\eta,\mathrm{i}}=\exp(-\mathrm{i}\eta(2X+Y)/\sqrt{5})$.
  3. 3.
    “random”: The initial state error is drawn randomly in each experiment for each $\beta\in\{+,\mathrm{i}\}$ and $\omega$, namely, $E_{\eta}=\exp(-\mathrm{i}\eta K)$ and $K$ is a random normalized Hermitian error generator.
The numerical results are given in Figure 14. These results suggest that the error scaling depends weakly on the strategy of adding initial state error. The study in Figure 14(a) suggests that the estimation error is linearly additive in the initial state error $\eta$, which agrees with our perturbative analysis. Moreover, we see that when $\theta=10^{-3}$ and $\eta=10^{-3}$, the induced estimation error of $\theta$ is around $10^{-6}$ and that of $\varphi$ is around $10^{-3}$ which are negligible compared to their own magnitudes. It is also worth noting that the plateau around $10^{-7}$ of $\theta$ estimation error (left panel in Figure 14(a)) is because the approximation error derived in Section Supplementary Information 3 dominates the estimation error. In Figure 14(b), we study the error scaling in a wide range of $\theta$ values. We see that the function form is nontrivial across different $\theta$ values. When $\theta$ is around $10^{-3}$, the $\theta$ estimation error scales quadratically in $\theta$, which might be due to some complex cancellation in Fourier transformation. The $\varphi$ estimation error scales reciprocally in $\theta$, because the signal-to-noise ratio of the complex phase increases as $\theta$ gets large. It agrees with the analysis in Section Supplementary Information 3. Moreover, Figure 14(c) shows that the spam error can be mitigated by increasing the circuit depth as the estimations error decays as $d^{-1}$ and $d^{-2}$ for $\theta$ and $\varphi$ respectively. Theoretically understanding these error scalings will be part of our future work.
(a) Estimation error as a function of initial state error $\eta$ ($\theta=10^{-3},d=20$)
(b) Estimation error as a function of swap angle $\theta$ ($\eta=10^{-6},d=20$)
(c) Estimation error as a function of depth parameter $d$ ($\theta=10^{-5},\eta=10^{-6}$)
Supplementary Figure 14: Estimation error induced by initial state error. We set $\varphi=3\pi/16$ and $\chi=5\pi/32$ for all experiments.
Supplementary Information 8 Computing the polynomial representation on a special set of points
@@
Furthermore, $Q_{\omega}^{(d)}$ and $\mathrm{Im}\left(e^{-\mathrm{i}\omega}P_{\omega}^{(d)}\right)$ can be determined from the recurrence relation in Equations 182 and 183
@@
Supplementary Information 9 Analysis of periodic calibration: variance lower bound and shortcomings
Supplementary Information 9.1 Overview of the methodology of periodic calibration
In this subsection, we provide an overview of periodic calibration, also known as Floquet calibration, which is proposed in [3, Appendix C] and [26, Appendix A]. Periodic calibration is a generalization of the robust single-qubit gate calibration [16] to entangling gates. The main component of the quantum circuit for performing periodic calibration is equivalent to the periodic part used in our method (see the shaded part in Figure 1 in the main text). However, instead of the initialization in terms of Bell states, periodic calibration measures the transition probability between tensor product states |$\ket{01}$ and |$\ket{10}$, namely, between logical quantum state |$\ket{0_{\ell}}$ and |$\ket{1_{\ell}}$. The exact parametric expression of this probability can be derived from the results presented in Section Supplementary Information 3, which is consistent with that in literature (upon convention difference):
@@
Supplementary Information 9.2 Optimality analysis using Fisher information and Cramér-Rao bound
@@
As we derived in Equation 145 in Section Supplementary Information 6.4, the quantum Fisher information that a quantum circuit using $\ell$ $U$-gates can maximally contribute is $\mathfrak{F}_{\max}=4M\ell^{2}$. Hence, the coincidence between the classical Fisher information with phase-matching condition and the maximal quantum Fisher information indicates that the optimality is attained at phase-matching because the Fisher information attains its maximum.
@@
Supplementary Figure 15: Cramér-Rao lower bound (CRLB) on the estimation variance of $\theta$ using periodic calibration. The number of measurement samples is set to $M=1\times 10^{5}$.
Supplementary Information 9.3 Violation of phase-matching condition implies exponentially worse estimation variance
To understand the importance of the phase-matching condition and the plateau of estimation variance, we provide an approximation to the Fisher information when the swap angle $\theta$ is small. For simplicity, we focus on the quantum circuit with a fixed depth parameter $d$ rather than a series of circuits with variable depths. It is referred to as $J_{\theta}$ below. We remark that the focus on the fixed-depth Fisher information does not lose generality due to the additivity of Fisher information.
@@
Supplementary Figure 16: Approximation analysis of Cramér-Rao lower bound (CRLB). (a) Fisher information of a fixed depth $J_{\theta}$ (blue solid line) and its approximation in Equation 195 (orange dashed line). The shaded area is the principal peak with |$\absolutevalue{\varphi}\leq\pi/d$. We set $d=10$ and $M=1$. (b) CRLB (scatters) and its approximation (dashed lines) based on Equation 195. We set $M=1\times 10^{5}$.
@@
Supplementary Information 9.4 Practical challenges due to the complex optimization landscape
@@
Supplementary Figure 17: Landscape analysis of periodic calibration. We set the exact value $\theta=1$ for all figures. (a) Optimization landscape of different $\varphi$ angles. We set the maximum depth $d=2^{1}0=1024$ and consider an ideal case where infinite many measurement samples are used $M=\infty$. (b) Optimization landscape of different maximum depth parameters $d$. We consider an ideal case where infinite many measurement samples are used $M=\infty$, and the phase-matching condition is satisfied $\varphi=0$. (c) Optimization landscape of different numbers of measurement samples $M$. We set $d=2^{8}=256$. We consider an ideal case where the phase-matching condition is satisfied $\varphi=0$.